OpenStudy (anonymous):

If mmc096-2.jpg = 5x – 8, mmc096-3.jpg = 3y, and mmc096-4.jpg = 132, find the values of x and y.

OpenStudy (anonymous):

mmc096 -2.jpg = 5x-8 mmc096 -3.jpg = 3y mmc096 -4.jpg = 132 ------------------------- rewrite equations: a - 2b = 5x - 8 a - 3b = 3y a - 4b = 132 then: -4b = 132 - a substitute: a - (132 -a ) = 3y a - 132 + a = 3y 132 = 3y y = 132/3 y = 44

OpenStudy (anonymous):

x 2 (2a -8b) = 10x -16 but a-4b = 132 or a = 132 +4b i.e 2[(132 - 4b)] -8b) = 10 x -16

OpenStudy (anonymous):

should be simple enough to solve from there